# Personalized Learning Path Planing through Goal-Driven Learner State Modeling

### v1

@@ Abstract.
Personalized Learning Path Planning (PLPP) aims to design adaptive learning paths that align with individual goals. While large language models (LLMs) show potential in personalizing learning experiences, existing approaches often lack mechanisms for goal-aligned planning. We introduce Pxplore, a novel framework for PLPP that integrates a reinforcement-based training paradigm and an LLM-driven educational architecture. We design a structured learner state model and an automated reward function that transforms abstract objectives into computable signals. We train the policy combining supervised fine-tuning (SFT) and Group Relative Policy Optimization (GRPO), and deploy it within a real-world learning platform. Extensive experiments validate Pxplore’s effectiveness in producing coherent, personalized, and goal-driven learning paths. We release our code and dataset to facilitate future research 111 https://github.com/Pxplore/pxplore-algo.
Personalized Learning Path Planing, Learner State Modeling, Reinforcement Learning, Group Relative Policy Optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Personalizationccs: Computing methodologies Reinforcement learningccs: Computing methodologies Sequential decision making
@@ 1. Introduction
Figure 1. Comparison of traditional and LLM-based learning path planning. Traditional methods produce static paths, while LLMs struggle to capture long-term progress.
Learning is a sequential and cumulative process that varies across learner. Personalized learning path planning (PLPP) for each learner has been a central pursuit in education  (Ng and Fung, 2024; Bernacki et al., 2021; Maghsudi et al., 2021; Tetzlaff et al., 2021). As shown in Figure 1, traditional approaches to PLPP (Tiwari, 2023; De la Vall and Araya, 2023), such as collaborative filtering and knowledge graphs, have advanced the organization of heterogeneous learning resources. However, their personalization remains constrained by predefined resources and static learner profiles, leading to fragmented and learning experiences (Ji et al., 2024; Li et al., 2023a; Abu-Rasheed et al., 2024; Yang et al., 2023). Large language models (LLMs), in contrast, overcome these limitations through scalable learner modeling and dynamic content adaptation (Wen et al., 2024; Hu et al., 2024a, b; Bonner et al., 2023). They integrates multi-dimensional learner data, adapts instructional materials, and even generates new content in a fine-grained, context-aware manner (Wang et al., 2025; Chen et al., 2024; Fitrianto et al., 2024; Li et al., 2023b; Mohsin et al., 2023). Yet, despite this potential, LLMs still struggle to model sustained learning progression, as they often lack mechanisms to represent long-term learner development or to incorporate feedback from authentic educational interactions (Jung, 2024; Zhao et al., 2024; Xia et al., 2019; Shou et al., 2020).
@@
$\bullet$ Understanding diverse learner states and adapting to heterogeneous learning resources. Learner states, encompassing cognitive levels, interest, and motivations, evolve continuously throughout the learning process (An et al., 2025). This requires adaptive systems that can infer and respond to these dynamic changes. However, existing approaches often lack explicit mechanisms to model such evolving learner states or to integrate diverse learning resources (Hu et al., 2024b; Moundridou et al., 2024), limiting their ability to construct and deliver coherent personalized learning experiences (Ikwuanusi et al., 2023; Bernacki et al., 2021; Maghsudi et al., 2021).
@@ 2. Related Work
Dimension Item Details
Evidence: [Turn 36: ”What are BP and PDP?”]
Confidence: 0.65  Status: [NOT_ALIGNED]
Evidence: [Turn 43: ”What is Monte Carlo Tree Search?”]
Confidence: 0.75  Status: [NOT_ALIGNED]
Evidence: [Turn 20: ”Skip this page, teacher.”]
Confidence: 0.80  Status: [NOT_ALIGNED]
Evidence: [Turn 25: ”The development of AI is changing… impact…”]
Confidence: 0.68  Status: [NOT_ALIGNED]
@@ 3.1.3. Automated Abstract Reward Function
# Learner States
# Session # Interaction # $O_{L}$ $O_{S}$ $M_{I}$ $M_{E}$
Train $250$ $11,328$ $273$ $325$ $275$ $288$
Test $50$ $3,256$ $53$ $76$ $63$ $62$
Total $300$ $14,584$ $326$ $401$ $338$ $350$
@@ 3.2.2. Post-Planning Architecture
Adaptive Narrative Generation. Instead of displaying the selected action, the system generates a narrative bridge connecting the new content to the learner’s ongoing context, as shown in Figure 4. Given the learner profile $p_{t}$, learning history $H_{t}=\{k_{1},\dots,k_{t}\}$, and selected action $k_{t+1}^{*}$, the delivery model $\mathcal{D}(\cdot)$ generates personalized instructions:
@@
$\bullet$ Expandable Architecture. Our post-planning architecture is designed as a modular and extensible component, allowing integration of additional generative services for richer instructional experiences including Retrieval-Augmented Generation (RAG) for in-depth explanations or Multimodal Generation, such as producing dynamic slides or visual aids aligned with prior content. As a proof of concept, we implement a dynamic slide generator that synthesizes new slides from selected actions while preserving stylistic and formatting consistency. This maintains visual and cognitive coherence across materials and reduces the cognitive load associated with inconsistent presentation formats. Overall, the extensible design demonstrates the Pxplore’s potential to evolve with diverse educational ecosystems. We leave the comprehensive exploration of such generative extensions to future work.
Figure 4. Example of Pxplore’s post-planning architecture that enables adaptive delivery of selected action.
4. Experiment
Table 3. Comparative performance of baseline models evaluated by average pedagogical alignment rate (%), reward computation ($\mathcal{R(\cdot)}$), and total component count within the defined Learner State Model. Pxplore consistently achieves the highest alignment and reward values across dimensions, demonstrating the effectiveness of the GRPO-based optimization.
@@ 4.2.1. Profiling Module
Experimental Setup. Using the curated dataset introduced in the previous section, we generate a structured learner profile for each learning session following the three-stage framework in Section 3.2.1. These profiles are compared against GPT-4o baseline, which recieves the same raw session logs but only a single instruction to generate a profile following the same output schema. Each profile is independently rated by two human annotators and three LLM-judges (Claude-3.7-Sonnet (Anthropic, 2025), GPT-4o, and Deepseek-R1 (DeepSeek-AI, 2025)) on a 5-point Likert scale, all blind to source. Evaluations are provided across four pedagogical dimensions:
(1) Interest Identification Accuracy (IIA): The correctness and specificity of inferred topic interests in the generated profile.
(2) The overall Profile Completeness (OPC): Coverage of core attributes, including engagement, cognition, affect, and interests.
(3) Actionability and Utility (AU): The practical usefulness of the generated profile for guiding downstream planning.
(4) Interpretability and Explainability (IE): Clarity of reasoning and strength of evidence linking learner behaviors to inferred states.
Results. To ensure reliability, we aggregate results from both human annotators and LLM-judges (Figure 5). Human annotators consistently rate Pxplore higher across all dimensions, with strong inter-annotator agreement (Pearson correlation $r>0.8$). Notably, Pxplore achieved the largest improvement in IE ($17.7\%$), followed by IIA ($6.26\%$) and AU ($6.63\%$), resulting in an overall increase of $7.88\%$. Both human and LLM-judges highlight Pxplore’s ability to construct explicit ”evidence chain”, for example, linking a learner’s review behavior directly to a detected misconception. This structured interpretability is essential for trustworthy learner profiling and aligns with established educational principles. Among LLM-judges, we observe different scoring patterns. Claude-3.7-Sonnet is the most conservative, while Deepseek-R1 shows highest alignment with human judgment, particularly in IE, likely due to its extended reasoning depth. Despite the differences, all evaluators confirm a consistent performance gap, validating the accuracy, completeness, and interpretability of Pxplore’s profiling framework.
@@ 4.2.2. Planning Module
Figure 7. Comparison of Pxplore and baselines on P@1 and NDCG@k metrics, indicating the alignment precision of top-ranked action with the expert-annotated ”best” choice, as well as the overall ranking quality across the candidate set.
@@ 4.2.3. Delivery Module
Experimental Setup. We use the dataset from prior evaluations and compare Pxplore’s adapted content against the original materials. Two human evaluators and three LLM-judges assess each version on a 5-point Likert scale across five pedagogical dimensions:
(1) Contextual Coherence (CC): How the content is logically structured and flows smoothly from the prior context.
(2) Personalization & Empathy (PE): Appropriateness and alignment with the learner’s current goals, knowledge state, and needs.
(3) Pedagogical Scaffolding (PS): Effectiveness of the content in providing structured guidance and support for learning.
(4) Conciseness & Efficiency (CE): Clarity, brevity, and effectiveness of the language in conveying key ideas.
(5) Motivational Appeal (MA): Capacity of the content to inspire and sustain learner motivation and engagement.
@@ 4.3. Real-World Experiment
To validate the practical effectiveness of Pxplore framework in authentic learning contexts, we integrate it into a deployed online learning platform and conducted a user study to examine its impact on both learning outcomes and learner experience.
@@
Procedure. Each learning session lasts approximately one hour, designed to simulate a realistic learning context:
(1) Pre-Test (10 mins): Participants complete a diagnostic test of 10 multiple-choice questions related to the initial lesson.
(2) Initial Learning Session (30 mins): Participant engage with one complete lesson on the system at their own pace.
(3) Advanced Learning Session (15 mins): After completing the initial lesson, participants proceed through the next lesson determined by the assigned planning method.
(4) Post-Test and Survey (10 mins): Participants complete a follow-up assessment and a perception survey. The full test questions and survey details are provided in Appendix C.
Metrics. We evaluate both learning outcomes and learner experience. Learning outcomes are measured by pre- and post-test score improvement across multiple-choice questions and short-answer questions, while learner experience is assessed using a 5-point Likert-scale across following dimensions:
(1) Helpfulness (Help.): Perceived usefulness of learning path;
(2) Diversity (Div.): Variety of content in topics and formats;
(3) Relevance (Rel): Alignment with prior learning content;
(4) Personalization (Pers.: Adaptation to individual preferences;
(5) Clarity (Clar.): Learning path coherence and logical flow;
(6) Motivation (Mot.): Stimulation of interest and engagement;
(7) Understanding (Und.): Depth of comprehension gained;
(8) Satisfaction (Sat.): The overall satisfaction score.
@@
Overall, the real-world experiment shows that while retrieval-based approach offer reasonable baseline performance, Pxplore demonstrates advantage in constructing cohesive, context-aware, and motivational learning paths. The integration of pre- and post-planning architecture further enables Pxplore to transform static content sequences into adaptive learning journeys that enhance both learner satisfaction and sustained educational impact.
Figure 8. Comparison of Pxplore and the retrieval-based baseline in terms of learning outcomes (pre- vs. post-test improvements) and learner experience across evaluated dimensions.
@@ A.2. Module Prompts
We further include the system prompts employed in each major module of our framework: learner profiling (Table 6), learning path planning (Table 7), and adaptive delivery (Table 5).
@@ Appendix B Hyperparameters
Table 8. Training setups and hyperparameters.
@@ Appendix C Real-World Experiment
We include supplementary materials for the real-world experiment to facilitate replication and further analysis. All participants completed pre- and post-tests, followed by a perception questionnaire evaluating their learning experience.
@@
Table 11 lists the learner experience questionnaire used for subjective evaluation. Participants rated each item on a 5-point Likert scale ranging from “Strongly Disagree” to “Strongly Agree.” The questionnaire covers eight pedagogical dimensions, helpfulness, diversity, relevance, personalization, clarity, motivation, understanding, and satisfaction, providing a comprehensive assessment of perceived adaptivity and engagement. All anonymized survey responses and test results are available in our public repository for reproducibility and future research.

### v2

# Personalized Learning Path Planning through Goal-Driven Learner State Modeling

@@ Abstract.
Personalized Learning Path Planning (PLPP) aims to design adaptive learning paths that align with individual goals. While large language models (LLMs) show potential in personalizing learning experiences, existing approaches often lack mechanisms for goal-aligned planning. We introduce Pxplore, a novel framework for PLPP that integrates a reinforcement-based training paradigm and an LLM-driven educational architecture. We design a structured learner state model and an automated reward function that transforms abstract objectives into computable signals. We train the policy combining supervised fine-tuning (SFT) and Group Relative Policy Optimization (GRPO), and deploy it within a real-world learning platform. Extensive experiments validate Pxplore’s effectiveness in producing coherent, personalized, and goal-driven learning paths. We release our code and dataset at  https://github.com/Pxplore/pxplore-algo.
Personalized Learning Path Planning; Learner State Modeling; Reinforcement Learning; Group Relative Policy Optimization
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792245submissionid: rfp1160ccs: Information systems Personalizationccs: Computing methodologies Reinforcement learningccs: Computing methodologies Sequential decision making
@@ 1. Introduction
Learning is a sequential and cumulative process that varies across learner. Personalized learning path planning (PLPP) for each learner has been a central pursuit in education  (Ng and Fung, 2024; Maghsudi et al., 2021). As shown in Figure 1, traditional approaches to PLPP (Tiwari, 2023; De la Vall and Araya, 2023), such as collaborative filtering and knowledge graphs, have advanced the organization of heterogeneous learning resources. However, their personalization remains constrained by predefined resources and static learner profiles, leading to fragmented and learning experiences (Ji et al., 2024; Abu-Rasheed et al., 2024; Li et al., 2023). Large language models (LLMs), in contrast, overcome these limitations through scalable learner modeling and dynamic content adaptation (Wen et al., 2024; Hu et al., 2024b, a). They integrates multi-dimensional learner data, adapts instructional materials, and even generates new content in a fine-grained, context-aware manner (Wang et al., 2025; Chen et al., 2024; Fitrianto et al., 2024). Yet, despite this potential, LLMs still struggle to model sustained learning progression, as they often lack mechanisms to represent long-term learner development or to incorporate feedback from authentic educational interactions (Jung, 2024; Zhao et al., 2024).
Figure 1. Traditional methods produce static paths, while LLMs struggle to capture long-term progress.
@@
$\bullet$ Understanding diverse learner states and adapting to heterogeneous learning resources. Learner states, evolving continuously throughout the learning process (An et al., 2025), requires adaptive systems that can infer and respond to these dynamic changes. However, existing approaches often lack explicit mechanisms to model such evolving learner states or to integrate diverse learning resources (Hu et al., 2024a; Moundridou et al., 2024), limiting their ability to construct and deliver coherent personalized learning experiences (Ikwuanusi et al., 2023; Bernacki et al., 2021; Maghsudi et al., 2021).
@@ 3.1.3. Automated Abstract Reward Function
Dimension Item Details
Evidence: [Turn 36: ”What are BP and PDP?”]
Confidence: 0.65  Status: [NOT_ALIGNED]
Evidence: [Turn 43: ”What is Monte Carlo Tree Search?”]
Confidence: 0.75  Status: [NOT_ALIGNED]
Evidence: [Turn 20: ”Skip this page, teacher.”]
Confidence: 0.80  Status: [NOT_ALIGNED]
Evidence: [Turn 25: ”The development of AI is changing… impact…”]
Confidence: 0.68  Status: [NOT_ALIGNED]
@@ 3.1.4. Policy Training
# Learner States
# Session # Interaction # $O_{L}$ $O_{S}$ $M_{I}$ $M_{E}$
Train $250$ $11,328$ $273$ $325$ $275$ $288$
Test $50$ $3,256$ $53$ $76$ $63$ $62$
Total $300$ $14,584$ $326$ $401$ $338$ $350$
@@ 3.2.2. Post-Planning Architecture
Adaptive Narrative Generation. Instead of displaying the selected action, the system generates a narrative bridge connecting the new content to the learner’s ongoing context. Given the learner profile $p_{t}$, learning history $H_{t}=\{k_{1},\dots,k_{t}\}$, and selected action $k_{t+1}^{*}$, the delivery model $\mathcal{D}(\cdot)$ generates personalized instructions:
@@
$\bullet$ Expandable Architecture. Our post-planning architecture is designed as a modular and extensible component, allowing integration of additional generative services for richer instructional experiences including Retrieval-Augmented Generation (RAG) for in-depth explanations or Multimodal Generation. As a proof of concept, we implement a dynamic slide generator that synthesizes new slides from selected actions while preserving stylistic and formatting consistency. This maintains visual and cognitive coherence across materials and reduces the cognitive load associated with inconsistent presentation formats. We leave the comprehensive exploration of such generative extensions to future work.
Table 3. Performance evaluated by average pedagogical alignment rate (%), reward computation ($\mathcal{R(\cdot)}$) within the defined Learner State Model. Pxplore consistently achieves the highest alignment rate and reward values.
@@ 4.2.1. Profiling Module
Experimental Setup. Using the curated dataset introduced in the previous section, we generate a structured learner profile for each learning session following the three-stage framework in Section 3.2.1. These profiles are compared against GPT-4o baseline, which recieves the same raw session logs but only a single instruction to generate a profile following the same output schema. Each profile is independently rated by two human annotators and three LLM-judges (Claude-3.7-Sonnet (Anthropic, 2025), GPT-4o, and Deepseek-R1 (DeepSeek-AI, 2025)) on a 5-point Likert scale, all blind to source. Evaluations are provided across four pedagogical dimensions: (1) Interest Identification Accuracy (IIA): The correctness and specificity of inferred topic interests in the generated profile. (2) The overall Profile Completeness (OPC): Coverage of core attributes, including engagement, cognition, affect, and interests. (3) Actionability and Utility (AU): The practical usefulness of the generated profile for guiding downstream planning. (4) Interpretability and Explainability (IE): Clarity of reasoning and strength of evidence linking learner behaviors to inferred states.
Results. To ensure reliability, we aggregate results from both human annotators and LLM-judges (Figure 4). Human annotators consistently rate Pxplore higher across all dimensions, with strong inter-annotator agreement (Pearson correlation $r>0.8$). Notably, Pxplore achieved the largest improvement in IE ($17.7\%$), followed by IIA ($6.26\%$) and AU ($6.63\%$), resulting in an overall increase of $7.88\%$. Both human and LLM-judges highlight Pxplore’s ability to construct explicit ”evidence chain”, for example, linking a learner’s review behavior directly to a detected misconception. This structured interpretability is essential for trustworthy learner profiling and aligns with established educational principles. Among LLM-judges, we observe different scoring patterns. Claude-3.7-Sonnet is the most conservative, while Deepseek-R1 shows highest alignment with human judgment. Despite the differences, all evaluators confirm a consistent performance gap, validating the accuracy, completeness, and interpretability of Pxplore’s profiling framework.
Figure 6. Results on the alignment precision of top-ranked action with the expert-annotated ”best” choice, as well as the overall ranking quality across the candidate set.
@@ 4.2.3. Delivery Module
Experimental Setup. We use the dataset from prior evaluations and compare Pxplore’s adapted content against the original materials. Two human evaluators and three LLM-judges assess each version on a 5-point Likert scale across five pedagogical dimensions: (1) Contextual Coherence (CC): How the content is logically structured and flows smoothly from the prior context. (2) Personalization & Empathy (PE): Appropriateness and alignment with the learner’s current goals, knowledge state, and needs. (3) Pedagogical Scaffolding (PS): Effectiveness of the content in providing structured guidance and support for learning. (4) Conciseness & Efficiency (CE): Clarity, brevity, and effectiveness of the language in conveying key ideas. (5) Motivational Appeal (MA): Capacity to inspire and sustain learner motivation and engagement.
@@ 4.3. Real-World Experiment
To validate the practical effectiveness of Pxplore framework in authentic learning contexts, we integrate it into a deployed online learning platform and conducted a controlled user study.
@@
Procedure. Each learning session lasts approximately one hour, designed to simulate a realistic learning context: (1) Pre-Test (10 mins): Participants complete a diagnostic test of 10 multiple-choice questions related to the initial lesson. (2) Initial Learning Session (30 mins): Participant engage with one complete lesson on the system at their own pace. (3) Advanced Learning Session (15 mins): After completing the initial lesson, participants proceed through the next lesson determined by the assigned planning method. (4) Post-Test and Survey (10 mins): Participants complete a follow-up assessment and a perception survey.
Metrics. We evaluate both learning outcomes and learner experience. Learning outcomes are measured by pre- and post-test score improvement across multiple-choice questions and short-answer questions, while learner experience is assessed using a 5-point Likert-scale across following dimensions: (1) Helpfulness (Help.): Perceived usefulness of learning path; (2) Diversity (Div.): Variety in topics and formats; (3) Relevance (Rel): Alignment with prior learning content; (4) Personalization (Pers.: Adaptation to individual preferences; (5) Clarity (Clar.): Learning path coherence and logical flow; (6) Motivation (Mot.): Stimulation of interest and engagement; (7) Understanding (Und.): Depth of comprehension gained; (8) Satisfaction (Sat.): The overall satisfaction score.
@@
Figure 7. Comparison of Pxplore and retrieval-based baseline in terms of learning outcomes and learner experience.
@@ 5. Conclusion
Acknowledgements.
This work is supported by National Natural Science Foundation of China (No.62277033) and National Engineering Laboratory for Cyberlearning and Intelligent Technology. We also thank the Deng Feng Fund of Tsinghua University for supporting this research.
@@ A.2. Module Prompts
We further include the system prompts employed in each major module of our framework: learner profiling (Table 6), learning path planning (Table 7), and adaptive delivery (Table 5). Figure 8 shows an example of content generated by the adaptive delivery module.
@@
Figure 8. Example of Pxplore’s post-planning architecture that enables adaptive delivery of selected action, connecting the new content to the learner’s on-going context.
@@ Appendix B Hyperparameters
Table 8. Training setups and hyperparameters used in our two-stage training paradigm.
@@ Appendix C Real-World Experiment
We include supplementary materials for the real-world experiment to facilitate the replication of our study and enable further analysis. To ensure a thorough evaluation of the learning outcomes, all participants completed pre- and post-tests, followed by a detailed perception questionnaire aimed at assessing their overall learning experience and engagement with the intervention.
@@
Table 11 lists the learner experience questionnaire used for subjective evaluation. Participants rated each item on a 5-point Likert scale ranging from “Strongly Disagree” to “Strongly Agree.” The questionnaire covers eight pedagogical dimensions, helpfulness, diversity, relevance, personalization, clarity, motivation, understanding, and satisfaction, providing a comprehensive assessment of perceived adaptivity and engagement.
To ensure transparency and facilitate future research, all anonymized survey responses and test results are made publicly available in our repository:  https://github.com/Pxplore/pxplore-algo.